# A polynomial bound on the pathwidth of graphs edge-coverable by $k$ shortest paths

### v1

@@ 1 Introduction
Graph covering is an old and recurrent topic in graph theory. For a fixed class $\mathcal{H}$ of graphs, the covering literature aims to answer two types of problems. First are optimization problems: given a graph $G$, what is the minimum number of graphs from $\mathcal{H}$ needed to “cover” [^1] Various notions of coverings can usually be considered here. $G$? We refer to [KU16] for a more general overview on the topic. The second type are structural problems: for the fixed class $\mathcal{H}$ of graphs and some $k\geq 1$, what is the structure of the graphs that can be “covered” by (at most) $k$ graphs from $\mathcal{H}$?
We say that a graph $G$ is edge-coverable (resp. vertex-coverable) by some subgraphs $H_{1},\ldots,H_{k}$ if every edge (resp. vertex) of $G$ belongs to at least one of the subgraphs $H_{i}$. Motivated by some algorithmic applications to the Isometric Path Cover problem and its variants, Dumas, Foucaud, Perez and Todinca [DFPT24] recently worked on a particular metric version of such covering problems. They proved that every graph edge-coverable (resp. vertex-coverable) by few shortest paths shares structural similarities with a simple path. More precisely, if a graph is edge-coverable or vertex-coverable by $k$ shortest paths, its pathwidth is smaller than an exponential in $k$ (see Section˜2 for a definition of pathwidth).
Theorem 1.1 ([DFPT24]).
@@
The proof of Theorem˜1.1 uses a branching approach to show that in every graph vertex-coverable (resp. edge-coverable) by at most $k$ shortest paths, for every vertex $u$ and every $i\geq 1$, the number of vertices at distance exactly $i$ from $u$ in $G$ is $O(k\cdot 3^{k})$ (resp. $O(3^{k})$). We also mention that a “coarse” variant of Theorem˜1.1 has recently been proved in [HP25]. Then, the bags of the path-decomposition proving Theorem˜1.1 are the union of every two consecutive layers in a BFS-layering of $G$ (starting from an arbitrary vertex). Such a path-decomposition is clearly not optimal in general: if we let $S_{h}$ denote the star with $h$ leaves, then $\mathrm{pw}(S_{h})=1$, while such a decomposition will always have width at least $\lfloor\tfrac{h}{2}\rfloor$, whatever the choice of $u$ is. The authors of [DFPT24] asked if the bounds from Theorem˜1.1 can be made polynomial in $k$. Our main result is a positive answer in the edge-coverable case.
@@
Our proof uses a different approach from [DFPT24]. Roughly speaking, it consists in finding some separator $X$ of polynomial size, such that for every component $C$ of GXG-X, there exists one path $P$ from the covering set of paths, such that $C$ has a “nice” layering with respect to $P$.
In terms of lower bounds, there exist for arbitrary values of $k\geq 1$ some graphs edge-coverable by $k$ shortest paths and with pathwidth equal to $k$ [DFPT24]. However, note that no better lower bound is known. Our second result is that this lower bound is tight for $k\leq 3$.
@@
Given a graph $G$, a subgraph $H$ of $G$ is isometric if every shortest path in $H$ is also a shortest path in $G$. For instance, shortest paths in $G$ are isometric subgraphs of $G$.
@@
Extending a result of Aigner and Fromme [AF84] for paths, Ball, Bell, Guzman, Hanson-Colvin and Schonsheck [BBG+17] proved that every graph vertex-coverable by $k$ isometric subtrees has cop number at most $k$. In particular, it is well-known that the cop number of a graph is always upper-bounded by its treewidth (see Section˜5 for a definition of treewidth). In analogy to Theorem˜1.1, a natural question is then the following: is it true that every graph which is vertex/edge-coverable by a small number of isometric subtrees has small treewidth? This question turns out to have a negative answer when considering vertex-coverability, as shown by the following example, pointed to us by Marcin Briański (personal communication): for every $t\geq 1$, let $G_{t}$ be the graph on $2t+2$ vertices obtained from the biclique $K_{t,t}$ after adding two vertices $a,b$ of degree $t$ respectively adjacent to all vertices from each of the sides of $K_{t,t}$ (see Figure˜1 for the case $t=3$). Then, $G_{t}$ is vertex-coverable by the two induced stars $S_{a},S_{b}$ respectively centered in $a$ and $b$ (which are in particular isometric subgraphs of $G_{t}$), but has treewidth at least $t$, as it contains $K_{t,t}$ as a minor. Our third result implies that such a construction does not exist anymore when considering graphs which are edge-coverable by two isometric trees.
@@
Routing problems play a central role not only in algorithmic graph theory and complexity, but also in structural graph theory. One of the best examples is the Disjoint Paths problem, appearing in the seminal Graph Minor series of papers of Robertson and Seymour [RS95], in which one is given as an input a graph $G$, an integer $k$ and $k$ pairs of vertices, and one must decide whether there exists $k$ pairwise disjoint paths connecting all pairs of vertices. This problem has been shown to be solvable in FPT time, when parameterized by the number $k$ of shortest paths (see [RS95, KKR12]), and this result has been a key ingredient to design efficient algorithms deciding the existence of minors in graphs. The metric variant Disjoint Shortest Path of this problem introduced in [ET98] (in which one further wants to decide whether there exist $k$ pairwise disjoint shortest paths connecting all input pairs of vertices) has attired a lot of attention, and was recently proved to be solvable in XP time, when parameterized by $k$ [BK17, Loc21, BNRZ23]. [^2] It is worth mentioning that all the aforementioned problems are NP-hard [FHW80, Loc21].
A number of metric variants of routing problems of similar flavour has been studied over the last decades. Among them, Isometric Path Cover studied in [AF84, FF01] takes as input a graph $G$ and an integer $k$, and asks whether $G$ is vertex-coverable by $k$-shortest paths. Its routing version Isometric Path Cover With Terminals takes as input a graph $G$, some integer $k$, and $k$ pairs of vertices, and asks whether there exist $k$ shortest paths that vertex-cover $G$ and connect all the input pairs of vertices. Both problems are NP-Complete [CDD+22, DFPT24]. Combining Theorem˜1.1 with Courcelle’s theorem [Cou90], the authors from [DFPT24] proved the existence of some algorithms running in XP and FPT time when parameterized by $k$, to solve respectively Isometric Path Cover and Isometric Path Cover With Terminals. In particular, it is still open whether Isometric Path Cover problem admits an FPT algorithm parameterized by $k$ [DFPT24]. We also refer to [CCFV26] for more results on Isometric Path Cover, and to [CDF+25] for some recent work on similar metric versions of routing problems.
@@
A number of metric variants of routing problems of similar flavour has been studied over the last decades. Among them, Isometric Path Cover studied in [AF84, FF01] takes as input a graph $G$ and an integer $k$, and asks whether $G$ is vertex-coverable by $k$-shortest paths. Its routing version Isometric Path Cover With Terminals takes as input a graph $G$, some integer $k$, and $k$ pairs of vertices, and asks whether there exist $k$ shortest paths that vertex-cover $G$ and connect all the input pairs of vertices. Both problems are NP-Complete [CDD+22, DFPT24]. Combining Theorem˜1.1 with Courcelle’s theorem [Cou90], the authors from [DFPT24] proved respectively the existence of some algorithms running in XP and FPT time when parameterized by $k$, to solve respectively Isometric Path Cover and Isometric Path Cover With Terminals. In particular, it is still open whether Isometric Path Cover problem admits an FPT algorithm parameterized by $k$ [DFPT24]. We also refer to [CDF+25] for some recent work on similar metric versions of routing problems. It is worth mentioning that Isometric Path Cover has be shown to be solvable in FPT time when parameterized by the isometric path complexity of a graph [CCFV26], a recent metric parameter of graphs introduced in [CDD+22] (under the name of isometric path antichain cover number).
@@
As mentioned above, the original parameterized algorithms from [DFPT24] solving Isometric Path Cover and its variant with terminals rely on Courcelle’s theorem, and thus the dependencies in $k$ are highly non-trivial and do not give rise in general to practical algorithms. In [CDF+25, Section 3], the authors designed a dynamic programming procedure solving the partition variant Isometric Path Partition of Isometric Path Cover in time $n^{O(\mathrm{tw}(G))}$, where $\mathrm{tw}(G)$ denotes the treewidth of the input graph. As pointed to us by Florent Foucaud (personal communication), it is very likely that combining our bound from Theorem˜1.2 with such an approach could lead to better running times than the ones given by Courcelle’s theorem for algorithms solving Isometric Path Edge-Cover and its version with terminals (the natural edge-covering variant of Isometric Path Cover).
@@
Section˜2 contains all preliminary definitions and some basic results related to pathwidth. Section˜3 contains a proof of Theorem˜1.2, Section˜4 contains a proof of Theorem˜1.3, and Section˜5 contains a proof of Theorem˜1.4. Note that Sections˜3, 4 and 5 are completely independent from each other. Section˜6 contains some discussion and additional questions left open by this work.
@@ 2.1 Notations and Basic Definitions
In this work, we consider undirected finite simple graphs, without loops. For a graph $G$, we denote with $V(G)$ its set of vertices, and $E(G)$ its set of edges. For any set $X\subseteq V(G)$ of vertices, $G[X]$ denotes the subgraph of $G$ induced by $X$, that is, the vertex set of $G[X]$ is $X$, and $G[X]$ contains all edges of $G$ with both end-points in $X$. GXG-X denotes the graph induced by $G$ on $V(G)\setminus X$. For simplicity, for every vertex $x$, we let GxG-x denote the graph $G-\{x\}$. For every subsets $X,Y\subseteq V(G)$ of vertices, we denote $E_{G}[X,Y]:=\{xy\in E(G):x\in X\setminus Y,y\in Y\setminus X\}$. Let $G^{\prime}$ be a graph, whose vertex set potentially intersects the one of $G$. The union of $G$ and $G^{\prime}$ is the graph $G\cup G^{\prime}:=(V(G)\cup V(G^{\prime}),E(G)\cup E(G^{\prime}))$.
For every $v\in V(G)$, we let $N_{G}(v)$ (or simply $N(v)$ when the context is clear) denote the set of neighbors of $v$ in $G$. Similarly, for every subset $X\subseteq V(G)$ of vertices, we let $N_{G}(X)$ (or simply $N(X)$) denote the set of vertices in $V(G)\setminus X$ having at least one neighbor in $X$. If $U\subseteq V(G)$, we similarly let $N_{U}(X)$ denote the set of vertices in $U\setminus X$ having at least one neighbor in $X$. For every $v\in V(G)$, we let $N_{G}(v)$ (or simply $N(v)$ when the context is clear) denote the set of neighbors of $v$. Similarly, for every subset $X\subseteq V(G)$ of vertices, we let $N_{G}(X)$ (or simply $N(X)$) denote the set of vertices in $V(G)\setminus X$ having at least one neighbor in $X$. If $U\subseteq V(G)$, we similarly let $N_{U}(X)$ denote the set of vertices in $U\setminus X$ having at least one neighbor in $X$.
@@
A path $P$ is a sequence of pairwise distinct vertices $v_{1},\ldots,v_{k}$ such that for each $i<k$, $v_{i}v_{i+1}\in E(G)$. A path connecting two vertices $v_{1}=x,v_{k}=y$ will also be called an $xy$-path. We call the vertices $v_{1},v_{k}$ the endvertices of $P$, while all its other vertices are its internal vertices. The length of $P$ corresponds to its number of edges. For each $i\leq j$, we let $P[v_{i},v_{j}]$ denote the subpath $v_{i},v_{i+1},\ldots,v_{j}$ of $P$, and we let $P^{-1}$ denote the path $v_{k},\ldots,v_{1}$. For every two paths $P=v_{1},\ldots,v_{p}$ and $Q=w_{1},\ldots,w_{q}$ such that $v_{p}=w_{1}$, and such that $P-v_{p}$ is disjoint from $Q-w_{1}$, we define the path $P\cdot Q:=v_{1},\ldots,v_{p}=w_{1},\ldots,w_{q}$. If $P=v_{1},\ldots,v_{k}$ is a path, for every $1\leq i<j\leq k$, we say that $v_{i}$ is before $v_{j}$ on $P$, and that $v_{j}$ is after $v_{i}$ on $P$. Note that by definition, a path is a sequence of vertices; however, by abuse of notations, we will often amalgamate a path $P=v_{1},\ldots,v_{k}$ of $G$ with its corresponding subgraph in $G$ (that is, the graph with vertices $v_{1},\ldots,v_{k}$ and edges $v_{i}v_{i+1}$ for each $i$). A path $P$ is vertical with respect to a vertex $u$, if for every $i\in\mathbb{N}$, $P$ has at most one vertex at distance $i$ from $u$ in $G$. Note that every path vertical with respect to some vertex $u$ is a shortest path, and that every shortest path starting from $u$ is vertical with respect to $u$.
@@
Let $G$ be a graph and $H_{1},\ldots,H_{k}$ be subgraphs of $G$. We say that the graphs $H_{1},\ldots,H_{k}$ edge-cover $G$ if $E(G)=\bigcup_{i=1}^{k}E(H_{i})$. Similarly, we say that $H_{1},\ldots,H_{k}$ vertex-cover $G$ if $V(G)=\bigcup_{i=1}^{k}V(H_{i})$.
@@
We let $d_{G}(\cdot,\cdot)$ (or simply $d(\cdot,\cdot)$ when the context is clear) denote the shortest-path metric of a graph $G$. Recall that a graph $H$ is an isometric subgraph of $G$ if $H$ is a subgraph of $G$ such that every shortest path in $H$ is also a shortest path in $G$. For every vertex $v$ of a graph $G$, we let $\mathrm{ecc}(v):=\max\{d_{G}(u,v):u\in V(G)\}\in\mathbb{N}\cup\{\infty\}$ denote the eccentricity of $v$.

### v2

@@ 1 Introduction
Graph covering is an old and recurrent topic in graph theory. For a fixed class $\mathcal{H}$ of graphs, the covering literature aims to answer two types of problems. First are optimization problems: given a graph $G$, what is the minimum number of graphs from $\mathcal{H}$ needed to “cover” [^1] Various notions of coverings can usually be considered here. $G$? We refer to Knauer and Ueckerdt [KU16] for a more general overview on the topic. The second type are structural problems: for the fixed class $\mathcal{H}$ of graphs and some $k\geq 1$, what is the structure of the graphs that can be “covered” by (at most) $k$ graphs from $\mathcal{H}$?
We say that a graph $G$ is edge-coverable (resp. vertex-coverable) by some subgraphs $H_{1},\ldots,H_{k}$ if every edge (resp. vertex) of $G$ belongs to at least one of the subgraphs $H_{i}$. Motivated by some algorithmic applications to the Isometric Path Cover problem and its variants, Dumas, Foucaud, Perez and Todinca [DFP+24] recently worked on a particular metric version of such a covering problem. They proved that every graph edge-coverable (resp. vertex-coverable) by few shortest paths shares structural similarities with a simple path. More precisely, if a graph is edge-coverable or vertex-coverable by $k$ shortest paths, its pathwidth is bounded by an exponential in $k$ (see Section˜2 for a definition of pathwidth).
Theorem 1.1 ([DFP+24]).
@@
The proof of Theorem˜1.1 uses a branching approach to show that, in every graph vertex-coverable (resp. edge-coverable) by at most $k$ shortest paths, and for every vertex $u$ and every $i\geq 1$, the number of vertices at distance exactly $i$ from $u$ in $G$ is $O(k\cdot 3^{k})$ (resp. $O(3^{k})$). We also mention that a “coarse” variant of Theorem˜1.1 has recently been proved [HP25]. Then, the bags of the path decomposition proving Theorem˜1.1 are the union of every two consecutive layers in a BFS-layering of $G$ (starting from an arbitrary vertex). Such a path decomposition is clearly not optimal in general: if we let $S_{h}$ denote the star with $h$ leaves, then $\mathrm{pw}(S_{h})=1$, while such a decomposition will always have width at least $\lfloor\tfrac{h}{2}\rfloor$, whatever the choice of $u$ is. Dumas et al. [DFP+24] asked if the bounds from Theorem˜1.1 can be made polynomial in $k$. Our main result is a positive answer in the edge-coverable case.
@@
Our proof uses a different approach from the one of [DFP+24]. Roughly speaking, it consists in finding some separator $X$ of polynomial size, such that for every component $C$ of GXG-X, $C$ has a “nice” layering with respect to one of the $k$ covering paths.
In terms of lower bounds, we provide for every $k\geq 1$, a construction of a graph edge-coverable by $k$ shortest paths and with pathwidth equal to $k$.
Proposition 1.3.
For any integer $k\geq 1$, there exists a graph $G_{k}$ that is
  1. (a)
    edge-coverable by $k$ shortest paths, and
  2. (b)
    that has pathwidth at least $k$.
To our knowledge, no better lower bound is known. Our second main result is that this lower bound is tight for $k\leq 3$.
@@
Given a graph $G$, a subgraph $H$ of $G$ is isometric if for every two vertices $x,y$ of $H$, the distances between $x$ and $y$ are the same in $G$ and $H$. In other words, if $x$ and $y$ are connected in $G$, then there must exist a shortest path between them in $G$ which is included in $H$.
@@
Extending a result of Aigner and Fromme [AF84] for paths, Ball, Bell, Guzman, Hanson-Colvin and Schonsheck [BBG+17] proved that every graph vertex-coverable by $k$ isometric subtrees has cop number at most $k$. In particular, it is well-known that the cop number of a graph is always upper-bounded by its treewidth (see Section˜5 for a definition of treewidth). In analogy to Theorem˜1.1, a natural question is then the following: is it true that every graph which is vertex/edge-coverable by a small number of isometric subtrees has small treewidth? This question turns out to have a negative answer when considering vertex-coverability, as shown by the following example, pointed to us by Marcin Briański (personal communication): for every $t\geq 1$, let $G_{t}$ be the graph on $2t+2$ vertices obtained from the biclique $K_{t,t}$ after adding two vertices $a,b$ of degree $t$ respectively adjacent to all vertices from each of the sides of $K_{t,t}$ (see Figure˜1 for the case $t=3$). Then, $G_{t}$ is vertex-coverable by the two induced stars $S_{a},S_{b}$ respectively centered in $a$ and $b$ (which are in particular isometric subgraphs of $G_{t}$), but has treewidth at least $t$, as it contains $K_{t,t}$ as a minor. Our third result implies that any such construction fails when considering graphs which are edge-coverable by two isometric trees, as we show that they have treewidth at most $2$.
@@
Routing problems play a central role not only in algorithmic graph theory and complexity, but also in structural graph theory. One of the best examples is the Disjoint Paths problem, appearing in the seminal Graph Minor series of papers of Robertson and Seymour [RS95], in which one is given as an input a graph $G$, an integer $k$ and $k$ pairs of vertices, and one must decide whether there exists $k$ pairwise disjoint paths connecting all pairs of vertices. This problem has been shown to be solvable in FPT time, when parameterized by the number $k$ of shortest paths [RS95, KKR12], and this result has been a key ingredient to design efficient algorithms deciding the existence of minors in graphs. The metric variant Disjoint Shortest Path of this problem [EIL98] (in which one further wants to decide whether there exist $k$ pairwise disjoint shortest paths connecting all input pairs of vertices) has attracted significant attention, and was recently proved to be solvable in XP time, when parameterized by $k$ [BK17, LOC21, BNR+23]. [^2] It is worth mentioning that all the aforementioned problems are NP-hard [FHW80, LOC21].
@@
A number of metric variants of routing problems of similar flavor has been studied over the last decades. Among them, Isometric Path Cover [AF84, FF01] takes as input a graph $G$ and an integer $k$, and asks whether $G$ is vertex-coverable by $k$ shortest paths. Its routing version Isometric Path Cover With Terminals takes as input a graph $G$, some integer $k$, and $k$ pairs of vertices, and asks whether there exist $k$ shortest paths that vertex-cover $G$ and connect all the input pairs of vertices. Both problems are NP-complete [CDD+22, DFP+24]. By combining Theorem˜1.1 with Courcelle’s theorem [COU90], it was shown [DFP+24] that there exists algorithms running in XP (respectively FPT time) when parameterized by $k$, to solve Isometric Path Cover (respectively Isometric Path Cover With Terminals). In particular, it is still open [DFP+24] whether the Isometric Path Cover problem admits an FPT algorithm parameterized by $k$. We also refer to Chakraborty et al. [CDF+25] for some recent work on similar metric versions of routing problems. It is worth mentioning that Isometric Path Cover admits some polynomial time $c$-approximation algorithms on graphs with isometric path complexity at most $c$ [CCF+26], a recent metric parameter of graphs [CDD+22] (introduced under the name of isometric path antichain cover number).
@@
As mentioned above, the original parameterized algorithms from [DFP+24] solving Isometric Path Cover and its variant with terminals rely on Courcelle’s theorem, and thus the dependencies in $k$ are highly non-trivial and do not give rise in general to practical algorithms. Chakraborty et al. [CDF+25, Section 3] designed a dynamic programming procedure solving the partition variant Isometric Path Partition of Isometric Path Cover in time $n^{O(\mathrm{tw}(G))}$, where $\mathrm{tw}(G)$ denotes the treewidth of the input graph. As pointed to us by Florent Foucaud (personal communication), it is very likely that combining our bound from Theorem˜1.2 with such an approach could lead to better running times than the ones given by Courcelle’s theorem for algorithms solving Isometric Path Edge-Cover and its version with terminals (the natural edge-covering variant of Isometric Path Cover).
@@
Section˜2 contains all preliminary definitions and some basic results related to pathwidth. Sections 3, 4, 5, and 6 contain respectively proofs of Theorems 1.2, 1.3, 1.4, and 1.5. Note that these sections are completely independent of each other, and can be read in any order. Section˜7 contains some discussion and additional questions left open by this work.
@@ 2.1 Notations and Basic Definitions
In this work, we consider undirected finite simple graphs, without loops. For a graph $G$, we denote by $V(G)$ its set of vertices, and $E(G)$ its set of edges. For any set $X\subseteq V(G)$ of vertices, $G[X]$ denotes the subgraph of $G$ induced by $X$, that is, the graph with vertex set $X$, and containing all edges of $G$ with both endpoints in $X$. GXG-X denotes the subgraph of $G$ induced by $V(G)\setminus X$. For simplicity, for every vertex $x$, we let GxG-x denote the graph $G-\{x\}$. For every subsets $X,Y\subseteq V(G)$ of vertices, we denote $E_{G}[X,Y]:=\{xy\in E(G):x\in X\setminus Y,y\in Y\setminus X\}$. Let $G^{\prime}$ be a graph, whose vertex set potentially intersects the one of $G$. The union of $G$ and $G^{\prime}$ is the graph $G\cup G^{\prime}:=(V(G)\cup V(G^{\prime}),E(G)\cup E(G^{\prime}))$.
For every $v\in V(G)$, we let $N_{G}(v)$ (or simply $N(v)$ when the context is clear) denote the set of neighbors of $v$ in $G$. Similarly, for every subset $X\subseteq V(G)$ of vertices, we let $N_{G}(X)$ (or simply $N(X)$) denote the set of vertices in $V(G)\setminus X$ having at least one neighbor in $X$. If $U\subseteq V(G)$, we similarly let $N_{G,U}(X)$ denote the set of vertices in $U\setminus X$ having at least one neighbor in $X$. When the context is clear, we will simply write $N_{U}(X)$.
@@
A path $P$ is a sequence of pairwise distinct vertices $v_{1},\ldots,v_{k}$ such that for each $i<k$, $v_{i}v_{i+1}\in E(G)$. Note that by definition, a path is a sequence of vertices; however, by abuse of notations, we will often identify a path $P=v_{1},\ldots,v_{k}$ of $G$ and its corresponding subgraph in $G$ (that is, the graph with vertices $v_{1},\ldots,v_{k}$ and edges $v_{i}v_{i+1}$ for each $i$ such that $1\leq i<k$). A path connecting two vertices $v_{1}=x$ and $v_{k}=y$ will also be called an $xy$-path. We call the vertices $v_{1}$ and $v_{k}$ the endvertices of $P$, while all its other vertices are its internal vertices. The length of $P$ corresponds to its number of edges. For each $i\leq j$, we let $P[v_{i},v_{j}]$ denote the subpath $v_{i},v_{i+1},\ldots,v_{j}$ of $P$, and we let $P^{-1}$ denote the path $v_{k},\ldots,v_{1}$. For every two paths $P=v_{1},\ldots,v_{p}$ and $Q=w_{1},\ldots,w_{q}$ such that $v_{p}=w_{1}$, and such that $P-v_{p}$ is disjoint from $Q-w_{1}$, we define the path $P\cdot Q:=v_{1},\ldots,v_{p}=w_{1},\ldots,w_{q}$. If $P=v_{1},\ldots,v_{k}$ is a path, for every $1\leq i<j\leq k$, we say that $v_{i}$ is before $v_{j}$ on $P$, and that $v_{j}$ is after $v_{i}$ on $P$. A path $P$ is vertical with respect to a vertex $u$, if for every $i\in\mathbb{N}$, $P$ has at most one vertex at distance $i$ from $u$ in $G$. Note that every path vertical with respect to some vertex $u$ is a shortest path, and that every shortest path starting from $u$ is vertical with respect to $u$.
@@
Let $G$ be a graph and $H_{1},\ldots,H_{k}$ be subgraphs of $G$. We say that the graphs $H_{1},\ldots,H_{k}$ edge-cover $G$ if $E(G)=\bigcup_{i=1}^{k}E(H_{i})$. Similarly, we say that $H_{1},\ldots,H_{k}$ vertex-cover $G$ if $V(G)=\bigcup_{i=1}^{k}V(H_{i})$. Note that whenever $G$ does not contain an isolated vertex, then every edge-cover is also a vertex-cover.
@@
We let $d_{G}(\cdot,\cdot)$ (or simply $d(\cdot,\cdot)$ when the context is clear) denote the shortest-path metric of a graph $G$. Recall that a graph $H$ is an isometric subgraph of $G$ if $H$ is a subgraph of $G$ such that for every two vertices $x,y\in V(H)$, we have $d_{H}(x,y)=d_{G}(x,y)$. For every vertex $v$ of a graph $G$, we let $\mathrm{ecc}(v):=\max\{d_{G}(u,v):u\in V(G)\}\in\mathbb{N}\cup\{\infty\}$ denote the eccentricity of $v$.